A stone is dropped by a person from the top of a building which is 200m tall at the same time another stone was thrown upwards with a velocity of 50m/s by a person standing at the foot of the building find the time the stones meet


1
Expert's answer
2021-04-13T06:31:49-0400

"\text{Let n be a function showing the distance from}\newline
\text{ ground level to stone.}"

"\text{For a stone thrown from a building:}"

"h_1(t)= 200-\frac{gt^2}{2};"

"\text{for a stone thrown from the ground:}"

"h_2(t)=V_0t-\frac{gt^2}{2}= 50t -\frac{gt^2}{2}"

"\text{When stones meet:}"

"h_1(t) = h_2(t)"

"200-\frac{gt^2}{2}=50t-\frac{gt^2}{2};"

"t=4"

Answer:4 seconds meeting time of stones